OpenStudy (anonymous):

Find the range of the data set: 10, 18, 21, 28, 12, 20 A. 10 B. 12 C. 18 D. 20 How do you find the range?

OpenStudy (anonymous):

@HeatDude

OpenStudy (anonymous):

isnt the range the smallest number?

OpenStudy (anonymous):

A because range means subtract largest number from lowest number

OpenStudy (anonymous):

In that case it would be 20-10=10

OpenStudy (anonymous):

oh gotchya so it would be 18

OpenStudy (anonymous):

Nope. It would be A.10

OpenStudy (anonymous):

huh? 28 is the highest number?

OpenStudy (anonymous):

range is largest number subtracted to smallest number, here it would be 28 - 10 = 18

OpenStudy (anonymous):

Oops sorry yeah your right didn't see the 28

OpenStudy (anonymous):

yep! lol

OpenStudy (anonymous):

Haha sorry my bad

OpenStudy (anonymous):

so with this one is would be Find the range of the data set: -15, 11, 12, -10, 12, -4 A. 1 B. 3.5 C. 12 D. 27

OpenStudy (anonymous):

how would I go about doing this one?

OpenStudy (anonymous):

12 -15 = -3

OpenStudy (anonymous):

wouldnt it be 27?

OpenStudy (anonymous):

It would probably be 27 because 12( biggest number) subtract -4 smallest number = 27 but not sure DON'T COUNT ON MY VOTE!!!! Lol xD I would ask someone else

OpenStudy (anonymous):

wait why isn't 3 any of the choices?

OpenStudy (anonymous):

because 12 is the highest number and -15 is the lowest. if you subtract those you get 27.

OpenStudy (anonymous):

would you say thats correct guys?

OpenStudy (anonymous):

I'm not sure again

OpenStudy (anonymous):

no if you subtract 12 - 15 = -3?

OpenStudy (anonymous):

12- -15 theirs a negative in front of the 15

OpenStudy (anonymous):

lol srry im thinking way deeper into algebraxD

OpenStudy (anonymous):

Yes you guys are right it's 12 a(-15) = 27

OpenStudy (anonymous):

lol its okay:)

OpenStudy (anonymous):

Oops -*

OpenStudy (anonymous):

Hey can u help answer this for me2+3m=17 I have to solve using mental math

OpenStudy (anonymous):

2 + 3m = 17 -2 -2 3m = 15 m = 5

OpenStudy (anonymous):

Okay thanks:-) lol